OpenStudy (anonymous):

What is the geometric mean between 2 and 50?

OpenStudy (anonymous):

Geometric mean = (Max value + Min value) / 2 = (2+50) / 2 = ?

OpenStudy (phi):

see https://en.wikipedia.org/wiki/Geometric_mean the geometric mean of two numbers a and b is \[ \text{ geo. mean } = \sqrt{ab} \]

OpenStudy (amistre64):

arithmetic mean is half the sum of 2 values phi is correct for the geometric mean of 2 values